# ManiSkill3: GPU Parallelized Robotics Simulation and Rendering for Generalizable Embodied AI

### v1

@@ Acknowledgments
The authors sincerely thank members of the Hao Su lab and Hillbot for great discussion and feedback, including but not limited to Nicklas Hansen, Rokas Bendikas, Jiayuan Gu, James Hou. The authors also thank members of the open source community for spotting early bugs, various contributions, and providing feedback on the framework. We especially acknowledge the generous support from Qualcomm Embodied AI Fund and Hillbot Inc. Stone Tao is supported in part by the NSF Graduate Research Fellowship Program grant under grant No. DGE2038238.
Appendix A Environments and Robots List
@@
This section covers key implementation and optimization details of the general supported task categories. Moreover, comparisons are made when possible with other robot simulation frameworks that have similar tasks/categories. For a complete list with videos of every task, see ManiSkill3 Tasks Documentation. A sample of 16 of the supported robots, which include mobile manipulators, floating grippers, quadrupeds, etc. is displayed in Figure 11. For a complete list with details/pictures of every robot, see ManiSkill3 Robots Documentation.
A.1 Table Top Manipulation
Table-top manipulation is primarily related to controlling one or more robot arms to manipulate an object on a table. Robots like Franka Emika Panda and Universal Robots 5 fall under this category. Typical tasks may include picking up objects, inserting a peg, assembling a structure, pushing objects, etc.
Robosuite (Zhu et al., 2020) is an existing robotics framework with CPU simulation+rendering implementations of a few table top tasks. Isaac Lab (Mittal et al., 2023) has only one table-top manipulation task supported out of the box. ManiSkill3 has 10+ different table-top manipulation tasks, some of which leverage heterogeneous parallel simulation where each parallel simulated+rendered environment has completely different numbers of objects with different geometries.
@@ A.1 Table Top Manipulation
A.2 Mobile Manipulation
@@
Mobile manipulation here refers to tasks in which a robot arm has a mobile base. Robots like Fetch and Stretch fall into this category. Typical tasks may include placing objects on surfaces, opening cabinet doors/drawers, picking up objects off the ground, etc.
@@
Implementation Details: The default robot supported is Fetch. The mobile base in particular is not simulated by driving the wheels, and is modeled similarly to AI2-THOR (Kolve et al., 2017) and Habitat (Szot et al., 2021) with one joint controlling forward/backward movement and another controlling rotation of the base. The Fetch robot definition and collision meshes have further been tuned to be simpler for faster simulation. Several impossible self-collisions between some links have been explicitly ignored to speed up simulation.
A.3 Room Scale Environments
@@
ManiSkill3 provides out-of-the-box code to build the ReplicaCAD environment from Habitat (Szot et al., 2021) and all AI2-THOR environments (Kolve et al., 2017) using assets compiled by the authors of the Habitat Synthetic Scenes Dataset (Khanna et al., 2023). Photo-realism is also possible by turning on the ray-tracing shader options when creating the environment.
@@
AI2-THOR (Kolve et al., 2017) and Habitat (Szot et al., 2021) have long-horizon mobile manipulation tasks but rely on slow CPU simulation / slow rendering systems, and do not support contact-rich physics for manipulation (only magical grasp) and do not look photorealistic. Robocasa (Nasiriany et al., 2024) has contact-rich long-horizon mobile manipulation tasks in photorealistic room-scale environments. However, Robocasa simulates and renders these scenes at around 25FPS as it does not use GPU parallelized simulation and rendering. In contrast ManiSkill3 can simulate the complex ReplicaCAD environment up towards 2000+ FPS with rendering.
@@
Implementation Details: We further make several modifications to ReplicaCAD to make it completely interactive as some of the collision meshes for articulations were modelled incorrectly and thus did not support low-level grasping. Via CoACD (Wei et al., 2022) we run convex decomposition on objects in AI2-THOR scenes to generate simulatable non-convex collision meshes so those objects can e.g. be grasped and moved around correctly. Via manual annotation by ManiSkill3 authors, certain categories of objects in AI2-THOR are made to be kinematic so they cannot be moved around (e.g. tables, TVs, clocks, paintings) with the rest allowed to be dynamic to be fully simulated (e.g. apples, baseball bat, cups).
A.4 Locomotion
@@
Locomotion here refers to controlling robot joints to move a robot from one location to another. Quadrupeds such as AnyMAL-C and humanoids such as Unitree-H1 fall into this category. Typical tasks may include standing still, moving to a target location, running as fast as possible, etc.
This type of task can be found in Isaac Lab which has several predefined locomotion tasks with tuned rewards and sim2real results.
@@
Implementation Details: Similar to Isaac Lab, quadruped robots in locomotion tasks are modified such that the large majority of collision meshes are removed, leaving behind just the feet, ankles, and the body. Moreover following Isaac Lab, joint limits are significantly constrained such that random actions do not easily cause the robot to fall over.
A.5 Humanoid / Bi-manual Manipulation
@@
Tasks here refer to the use of a humanoid embodiment such as the Unitree H1 robot or bi-manual robot embodiments for manipulation tasks.
@@
A.6 Multi-Agent Robots
Multi-Agent robots refer to support for controlling multiple different robots in the same simulation to perform a task. Setups such as multiple quadrupeds or robot arms fall into this category. A common task is the handover of objects.
This type of task can be found in Robosuite. Isaac Lab supports this type of task, but not out of the box. Past environments based on older versions of Isaac have example tasks with multiple robot arms/hands running on GPU simulation (Chen et al., 2022).
Implementation Details: By default the action space is a dictionary action space in multi-agent environments with a dictionary key for each controllable agent. This action space can easily be flattened into a single vector if necessary via a environment wrapper.
A.7 Drawing/Cleaning
Drawing/Cleaning refers to tasks for dynamically ”adding/removing” objects to simulate the effect of drawing or cleaning. A task could be to draw the outline of a shape on a canvas or clean dirty spots on a table surface.
ManiSkill3 is the only framework that supports this kind of task with GPU parallelization.
@@ A.7 Drawing/Cleaning
A.8 Dextrous Manipulation
Dextrous Manipulation refers to tasks often involving multi-fingered hands and dense/rich contacts occurring during manipulation. An example task is in-hand rotation.
@@ A.8 Dextrous Manipulation
A.9 Vision Tactile Manipulation
Vision Tactile Manipulation refers to tasks that rely on processing tactile information like images to solve manipulation tasks. Tasks could include key insertion which require tactile inputs to solve due to visual occlusions.
ManiSkill3 is the only simulator with vision tactile manipulation tasks that also have sim2real capabilities.
Implementation Details: The vision-tactile, sim2real, manipulation environments are ported over from the 2024 ManiSkill Vision-based-Tactile Manipulation Skill Learning Challenge (Chen et al., 2024).
A.10 Classic Control
Classic control is quite broad but in this context refers to fake robots tasked with achieving some stable pose or moving in a direction at desired speeds. Tasks include cartpole balancing, humanoid walking, hopper etc.
DM-Control (Tunyasuvunakool et al., 2020) has the most implemented control tasks and Isaac Lab has a few GPU parallelized variants. ManiSkill3 has GPU parallelized simulation+rendering variants of the most control tasks.
Implementation Details: ManiSkill3 uses its MJCF loader to load the MJCF robot definitions from the original DM-Control repository and tunes robot pd joint delta position controllers to align as closely as possible to the behavior seen in DM-Control.
A.11 Digital Twins
@@
For real2sim environments, ManiSkill3 ports over and GPU parallelizes some environments from SIMPLER (Li et al., 2024), which enables efficient evaluation of policies trained on real world data like the generalist RT-1 and Octo models. The primary tricks include green-screening a real world image and texture matching which have been copied over and parallelized. In depth documentation/tutorials are provided on how to build custom digital twins for real2sim. We ensured the GPU parallelized port of SIMPLER achieves similar results/behaviors as the original CPU simulated environments as shown in Figure 12.
Figure 12: Evaluated success rates of generalist robotics models like Octo and RT-1X on 4 different tasks. The correlation and MMRV metrics are close to that of the original paper.
TRANSIC-Envs (Jiang et al., 2024) provide reproducible sim2real digital twin setups but rely on state estimation via ArUco markers for sim2real transfer and does not support visual feedback. Dextreme (Handa et al., 2023) provides a realistic in hand cube rotation environment but also does not support efficient visual feedback and relies on accurate state estimation. ManiSkill3 on the other hand has fast visual data generation that enable mimicing real-world cameras for training at scale.
A.12 Soft Body Manipulation Environments
@@
Appendix B Simulation and Rendering Benchmarking

### v2

@@ Acknowledgments
The authors sincerely thank members of the Hao Su lab and Hillbot for great discussion and feedback, including but not limited to Nicklas Hansen, Rokas Bendikas, Jiayuan Gu, James Hou. The authors also thank members of the open source community for bug finding, contributions, and feedback on this framework, including but not limited to Alexandre Brown, Gaotian Wang, Dwait Bhatt, Jeremy Morgan. We especially acknowledge the generous support from Qualcomm Embodied AI Fund and Hillbot Inc. Stone Tao is supported in part by the NSF Graduate Research Fellowship Program grant under grant No. DGE2038238.
Appendix Table of Contents
  • Environments and Robots VII
  • Reinforcement Learning Baselines IX
  • Imitation Learning Baselines X
  • Simulation and Rendering Benchmarking XI
  • VR Teleoperation XII
VII Environments and Robots List
@@
This section covers key implementation and optimization details of the general supported task categories. For a complete list with videos of almost every task, see ManiSkill3 Tasks Documentation. For videos of soft body and vision-tactile tasks we refer readers to the ManiSkill2 paper and A sample of 16 of the supported robots, which include mobile manipulators, floating grippers, quadrupeds, etc. is displayed in Figure 14. For a complete list with details/pictures of nearly every robot, see ManiSkill3 Robots Documentation.
We emphasize here that while ManiSkill3 may not have the most distinct number of tasks compared to some benchmarks, the core contribution is supporting a diverse array of possible tasks with open-sourced code that users can reference and use to easily build more tasks.
VII-A Table Top Manipulation
Figure 15: Example table-top manipulation task showing the start state and the solved state. The robot arm must grasp the white-orange peg and insert it into the box.
Table-top manipulation is primarily related to controlling one or more robot arms to manipulate an object on a table. Robots like Franka Emika Panda and Universal Robots 5 fall under this category. Typical tasks may include picking up objects, inserting a peg, assembling a structure, pushing objects, etc. An example task is shown in Figure 15.
@@ VII-A Table Top Manipulation
VII-B Mobile Manipulation
Figure 16: Example mobile manipulation task showing the start state and the solved state. The mobile robot arm must grasp the handle indicated by the green sphere and pull it open.
@@ VII-B Mobile Manipulation
Mobile manipulation here refers to tasks in which a robot arm has a mobile base. Robots like Fetch and Stretch fall into this category. Typical tasks may include placing objects on surfaces, opening cabinet doors/drawers, picking up objects off the ground, etc. Example task is shown in Figure 16.
@@
Implementation Details: The default robot supported is Fetch. The mobile base in particular is not simulated by driving the wheels, and is modeled similarly to AI2-THOR [29] and Habitat [46] with one joint controlling forward/backward movement and another controlling rotation of the base. The Fetch robot definition and collision meshes have further been tuned to be simpler for faster simulation. Several impossible self-collisions between some links have been explicitly ignored to speed up simulation.
VII-C Room Scale Environments
@@
ManiSkill3 provides out-of-the-box code to build the ReplicaCAD environment from Habitat [46], all AI2-THOR environments [29] using assets compiled by the authors of the Habitat Synthetic Scenes Dataset [27], and the RoboCasa scenes dataset [39]. Photo-realism is also possible by turning on the ray-tracing shader options when creating the environment.
@@
AI2-THOR [29] and Habitat [46] have long-horizon mobile manipulation tasks but rely on slow CPU simulation / slow rendering systems, and do not support contact-rich physics for manipulation (only magical grasp) and do not look photorealistic. Robocasa [39] has contact-rich long-horizon mobile manipulation tasks in photorealistic room-scale environments. However, Robocasa simulates and renders these scenes at around 25FPS as it does not use GPU parallelized simulation and rendering. In contrast ManiSkill3 can simulate the complex ReplicaCAD environment up towards 2000+ FPS with rendering.
@@
Implementation Details: We further make several modifications to ReplicaCAD to make it completely interactive as some of the collision meshes for articulations were modelled incorrectly and thus did not support low-level grasping. Via CoACD [51] we run convex decomposition on objects in AI2-THOR scenes to generate simulatable non-convex collision meshes so those objects can e.g. be grasped and moved around correctly. Via manual annotation by ManiSkill3 authors, certain categories of objects in AI2-THOR are made to be kinematic so they cannot be moved around (e.g. tables, TVs, clocks, paintings) with the rest allowed to be dynamic to be fully simulated (e.g. apples, baseball bat, cups) to optimize simulation speed.
VII-D Locomotion
Figure 17: Example locomotion task showing the start state and the solved state. The quadruped must use its 4 legs to to move to the green goal.
@@ VII-D Locomotion
Locomotion here refers to controlling robot joints to move a robot from one location to another. Quadrupeds such as AnyMAL-C and humanoids such as Unitree-H1 fall into this category. Example task is shown in Figure 17.
@@
Implementation Details: Similar to Isaac Lab, quadruped robots in locomotion tasks are modified such that the large majority of collision meshes are removed, leaving behind just the feet, ankles, and the body visualized in Figure 18. Moreover following Isaac Lab, joint limits are significantly constrained such that random actions do not easily cause the robot to fall over.
Figure 18: Comparison of the visual and collision mesh of one of the robot quadruped models, AnyMAL-C.
VII-E Humanoid / Bi-manual Manipulation
Figure 19: Example humanoid manipulation task showing the start state and the solved state. The humanoid must use both arms to grasp the box and transport it to the other table.
@@ VII-E Humanoid / Bi-manual Manipulation
Tasks here refer to the use of a humanoid embodiment such as the Unitree H1 robot or bi-manual robot embodiments for manipulation tasks. Example task shown in 19.
@@
VII-F Multi-Agent Robots
Figure 20: Example multi-agent robotics task showing the start state and solved state. The left robot must push a cube over to the other robot to pick up.
Multi-Agent robots refer to support for controlling multiple different robots in the same simulation to perform a task. Setups such as multiple quadrupeds or robot arms fall into this category. A common task is the handover of objects. Example task shown in Figure 20.
This type of task can be found in Robosuite. Isaac Lab supports this type of task, but not out of the box. Past environments based on older versions of Isaac have example tasks with multiple robot arms/hands running on GPU simulation [10].
Implementation Details: By default the action space is a dictionary action space in multi-agent environments with a dictionary key for each controllable agent, which follows the standard PettingZoo API [48]. PettingZoo is currently the most popular interface for multi-agent RL environments. For users who do not wish to do multi-agent RL they can flatten the action space into a single vector if necessary via a environment wrapper provided by ManiSkill3.
VII-G Drawing/Cleaning
Figure 21: Example drawing/cleaning task showing the start state and solved state. The robot must draw a given SVG file (which describes a set of lines), in this case the letter A.
Drawing/Cleaning refers to tasks for dynamically ”adding/removing” objects to simulate the effect of drawing or cleaning. A task could be to draw the outline of a shape on a canvas or clean dirty spots on a table surface. Example task shown in Figure 21.
ManiSkill3 is the only framework that supports this kind of task out of hte box with GPU parallelization and rendering.
@@ VII-G Drawing/Cleaning
VII-H Dextrous Manipulation
Figure 22: Example dextrous manipulation robotics task showing the start state and solved state. The robot hand must rotate the object in place to a desired orientation.
Dextrous Manipulation refers to tasks often involving multi-fingered hands and dense/rich contacts occurring during manipulation. An example task is in-hand rotation. Example task shown in Figure 22.
@@ VII-H Dextrous Manipulation
VII-I Vision Tactile Manipulation
Figure 23: Example task of a key insertion task in simulation (top left), the real world equivalent (top right), and plots of the tactile feedback data (bottom).
Vision Tactile Manipulation refers to tasks that rely on processing tactile information like images to solve manipulation tasks. Tasks could include key insertion which require tactile inputs to solve due to visual occlusions. Example task shown in Figure 23
Implementation Details: The vision-tactile, sim2real, manipulation environments are ported over from the 2024 ManiSkill Vision-based-Tactile Manipulation Skill Learning Challenge [9].
VII-J Classic Control
Figure 24: Example control environments: CartPoleBalance, Hopper, and Ant.
Classic control is quite broad but in this context refers to fake robots tasked with achieving some stable pose or moving in a direction at desired speeds. Tasks include cartpole balancing, hopper etc. Examples are shown in Figure 24
DM-Control [50] has the most implemented control tasks and Isaac Lab has a few GPU parallelized variants. ManiSkill3 has GPU parallelized simulation+rendering variants of the most control tasks.
Implementation Details: ManiSkill3 uses its MJCF loader to load the MJCF robot definitions from the original DM-Control repository and tunes robot pd joint delta position controllers to align as closely as possible to the behavior seen in DM-Control/Mujoco.
VII-K Digital Twins
@@
For real2sim, ManiSkill3 ports over and GPU parallelizes some environments from SIMPLER [31], which enables efficient evaluation of policies trained on real world data like the generalist RT-X and Octo models. The primary tricks include green-screening a real world image and texture matching which have been copied over and parallelized. We ensured the GPU parallelized port of SIMPLER achieves similar results/behaviors as the original CPU simulated environments as shown in Figure 25.
Figure 25: Evaluated success rates of generalist robotics models like Octo and RT-1X on 4 different tasks. The correlation and MMRV metrics are close to that of the original paper. MMRV is Mean Maximum Rank Violation (lower the better), which assess the accuracy of the consistency of the rankings of real and sim policy evaluations.
For sim2real, we provide tools useful for training policies in simulation and directly deploying into the real-world. Given the number of details involved in sim2real, we explain the implementation details of vision-based sim2real in detail in Appendix VIII.
VII-L Soft Body Manipulation Environments
@@
VIII Vision-Based Sim2Real
We describe how one can use ManiSkill3 to perform end-to-end vision-based sim2real by training with PPO a RGB-based manipulation policy in simulation and zero-shot deploying it in the real world. Past work such as TRANSIC-Envs [26] provide reproducible sim2real digital twin setups but rely on state estimation via ArUco markers for sim2real transfer and does not support visual feedback. Dextreme [21] provides a realistic in hand cube rotation environment but also does not support efficient visual feedback and relies on accurate state estimation. ManiSkill3 on the other hand has fast visual data generation that enable mimicing real-world cameras inputs for training at scale and fast. We note however the example setup here is a task less complex than that of e.g. Dextreme but point out that our reproducible setup is easily extendable and could be the start of opening a new avenue of sim2real research with efficient and fast large-scale simulation training on image inputs.
VIII-A Hardware
We use the low-cost \$300 Koch v1.1. robot arm and the LeRobot library for a simple accessible python interface to control the robot arm. Due to the simple but accessible hardware, it is generally harder to deploy policies on this robot as the motors are less precise compared to more expensive robot arms. Thus the vision-based sim2real demonstrated here can easily work on other robot arms. Apart from the robot we also setup one third view iPhone camera (any LeRobot compatible camera can work) to record RGB observation data.
VIII-B Controller / Action Space
We use a target delta joint position controller in both simulation and the real-world at a control frequency of 30Hz. At timestep $t=0$ after an (real or simulation) environment reset we save an initial target joint position value $\bar{q}_{t}=q_{t}$ where $q_{t}$ is the joint position of the robot arm at timestep $t$. Given an action $a_{t}$ at timestep $t$, we set $\bar{q}_{t+1}=\bar{q}_{t}+a_{t}$.
At each timestep $t$, we drive the robot’s joint positions to $\bar{q}_{t}$. In simulation this is equivalent to setting target joint position values in the PhysX physics engine and the behavior is defined by a PD controller. In the real world this is equivalent to setting the next joint position of the robot arm to be $\bar{q}_{t}$, and most robot arms have APIs that enable this functionality and will internally run their own control loop to reach the set joint position.
The actual action space of the environment is normalized to a range of -1 to 1 for each joint. Given a predicted action from a policy $\hat{a}_{t}$ we clip it first to the range of $[-1,1]$ and then unnormalize it to obtain $a_{t}=\frac{\hat{a}_{t}(h-l)}{2}$, which is the delta action applied to the $\bar{q}_{t}$. For the Koch robot arm $l=-0.05,h=0.05$, which represent a maximum change of 0.05 radians in either direction of each joint.
We choose the target delta joint position controller as the sim2real dynamics gap for quasi-static tasks (tasks where objects only move when the robot manipulate them) is minimized. In both simulation and real-world in quasi-static settings with small enough action sizes as long as within the timeframe of one control step the real world robot and simulated robot reach $\bar{q}_{t}$ at timestep $t$ then the simulation and real-world are aligned well. We note that we leave it to future work to explore controller and sim2real design for addressing more dynamic tasks such as catching a thrown object which require higher control frequencies.
VIII-C Observation Space
The observation space follows that of all visual RL baselines, which contains both proprioceptive data and the camera input RGB data. For the simulated tasks the prioprioceptive data contains both the current robot joint positions $q_{t}$ as well as the target joint positions $\bar{q}_{t}$ which is required to make the problem markov and more easily solvable via RL. The RGB data is a $128\times 128$ RGB image captured during each control step of the environment.
We further include a binary label of whether or not the robot is grasping something, which is a simple check of whether the joint position of the gripper joint is within 0.02 radians of the target joint position of the gripper joint. If the joint position of a robot cannot reach the target joint position this indicates something is blocking that joint, e.g. an object being grasped. Empirically we find that due to occlusions that occur once the cube is grasped due to the camera angle and the gripper link covering the cube, the trained policy has difficulty discerning whether or not it has grasped a cube. The binary is grasped feature generated from $q_{t},\bar{q}_{t}$ values is sufficient to overcome the issue of occlusions.
All of these observation data points are easily obtained in the real world without needing to set up perception stacks to observe e.g. object poses. No observation delays are added but can be useful for higher frequency tasks.
VIII-D Environment Setup and Domain Randomizations
Figure 26: Sample images of the start states of the simulated environment and real-world environments, showing the reset distribution. Videos of the reset distribution can be found on our website.
The simulated environment contains 4 objects, the floor, the table, the cube, and the robot arm. A visualization of the reset distribution can be found in Figure 26. We perform the following randomizations at each timestep of the environment:
  • Camera pose is randomized by randomly shifting the camera position up to 2.5 cm in x, y, and z axes in the world frame. This randomization scale sufficiently covers possible misalignments between simulated and real-world cameras.
We perform the following randomizations each time the environment is reset:
  • The initial robot joint positions $q_{0}$ are sampled from $\mathcal{N}(q_{R},0.02)$, where $q_{R}$ represents the joint positions we set for the real Koch v1.1 robot at each reset.
  • The position of the cube is uniformly sampled within a $10$cm2 square in front of the robot arm, and the cube’s z-axis rotation is randomized completely.
We perform the following randomizations each time the environment is reconfigured, which permit the environment to use different object assets, materials, camera settings. This occurs only during environment resets, but typically is done less often as it slows down reset times.
  • Cube side lengths are sampled uniformally in the range $[1.5,2.25]$ cm
  • Cube frictions are sampled from $\mathcal{N}(0.3,0.025)$
  • Cube color is sampled uniformally from all possible RGB values.
  • Robot chassis color is sampled from $\mathcal{N}(c,0.05)$, where $c$ is the base robot chassis RGB color scaled from 0 to 1.
  • Robot motor color is sampled from $\mathcal{N}(m,0.05)$, where $m$ is the base robot motor color scaled from 0 to 1.
VIII-E Reward
$\mathbb{I}_{grasped}$ indicates whether the robot is grasping the cube. $\mathbb{I}_{touch\_table}$ indicates if the robot is robot’s gripper is colliding with the table. $\mathbb{I}_{tcp\_close}$ indicates whether the tool center point, the point between the robot’s two gripper fingers $x^{tcp}_{t}$, is within a half of the cube’s side length from the cube’s position, $x^{cube}_{t}$. We also define the vector from the first gripper finger to the second as $\vec{g}$ and the cube size as $c$. The reward function, for a state $s$ and action $a$ is:
$r_{t}(s,a)=$ $\ r_{dist}(x^{tcp}_{t},x^{cube}_{t},15)+r_{orient}(\vec{g})$
$+2\mathbb{I}_{tcp\_close}r_{dist}(||\vec{g}||,c,40)$
$+\mathbb{I}_{grasped}\left(1+3r_{dist}(q_{t},q_{R},\frac{4}{\pi})\right)$
$-2\mathbb{I}_{touch\_table}$
$r_{dist}$ is a common distance reward function used in ManiSkill3, where $r_{dist}(x,y,z)=1-tanh(z*||x-y||)$. $r_{orient}(\vec{g})=1-(\vec{u}\cdot\frac{\vec{g}}{||\vec{g}||})$, where $\vec{u}$ is the up vector. Intuitively, this means rewarding the agent with a value of one for keeping $\vec{g}$ perpendicular to the table normal and a value of zero if it is parallel instead. $q_{t}$ and $q_{R}$ represent the joint positions at timestep $t$ and upon environment reset, respectively.
We find that action magnitude, $||a||$, does not need to be considered within the reward, as the controller’s restrictions of $l$ and $h$ are appropriately tuned for safe deployment and for the realistic maximum joint acceleration magnitudes for each real robots’ motors.
VIII-F Training
PPO with RGB inputs is used for training, details on the overall RL implementation is discussed in Appendix IX. Specific to this task we use a reconfiguration frequency of 10 to ensure the RL policy sees a sufficiently diverse set of randomizations over object geometry and materials while also ensuring we are not performing slow reconfiguration resets too often. Ultimately training takes about 1 hour on a 4090 GPU. The final training curves can be found in the main paper in Figure 13.
IX Reinforcement Learning Baselines
We most comprehensively test PPO and TD-MPC2 as reinforcement learning baselines, which represent two ongoing sides of RL research, walltime efficient RL and sample efficient RL. SAC is also provided but is not heavily benchmarked as of now. When possible RL baselines have sample-efficient and wall-time efficient hyperparameters. These baselines are implemented with PyTorch [1].
IX-A Proximal Policy Optimization (PPO)
Neural Network Architecture: For state based tasks the actor and critic networks are separated and are comprised of a MLP with three 256 unit hidden layers followed by Tanh activation. For vision based tasks the actor and critic share a feature processing backbone, namely a NatureCNN following the default setup that the popular Stable Baselines 3 RL library [42] uses.
Training Details and Hyperparameters: The key hyperparameters that were tuned per environment were largely the discount factor, GAE lambda factor, the max episode steps/horizon, as well as the number of parallel environments. Typically we try to use up to 4096 parallel environments when possible for state based training and about 256 to 1024 environments for RGB based training. Discount factor by default is kept at a low value of 0.8 which empirically works well for most tasks and increased up to 0.99 if 0.8 does not work. Episode horizon is tuned such that the task can be solved within that horizon and the horizon is not much longer than the optimal solve time. We further tune the number of steps sampled from each parallel environment in each rollout. Generally a value of 16 or 32 is used and is tuned down if possible. However, we observe most tasks cannot be solved with low number of steps sampled per parallel environment like PegInsertionSide, likely due to long horizon temporal dependencies that are harder to model.
For all experiments we always turn partial resets on, which enables environments to reset early upon reaching a fail or success state if there are any. Partial resets generally enable faster training by reducing the time spent sampling the environment in irrelevant states.
ManiSkill environments support setting environment reconfiguration freqencies, which control how often during an environment reset we will also permit resetting/randomizing simulation properties that are not changeable after starting simulation (such as the object shape as done in the PegInsertionSide task). By default all experiments have a reconfiguration frequency of 0 for training environments as generally with sufficiently high number of parallel environments there is enough diversity in the training data. For more advanced use-cases like training robust sim2real policies as done in Appendix VIII, we use a non-zero reconfiguration frequency to better randomize the object geometries and other properties.
Due to the number of hyperparameters and environments we do not include them in the paper itself and refer users to the public training runs on the Wandb page for PPO (state and RGB) results which show all metrics and configurations.
IX-B Temporal Difference Learning for Model Predictive Control 2 (TD-MPC2)
Neural Network Architecture: For tasks with state-based observations, we used the original TD-MPC2 architecture from the original paper. For vision-based tasks, we modified the encoder to handle larger 128×128 input images and additional external state information (e.g., joint position, goal position). In particular, we introduced an additional convolutional layer to the RGB encoder, as well as a separate MLP (composed of one 256-unit and one 512-unit layer-normed, linear layer) to process the extra state data. The outputs from the RGB encoder and state encoder are then combined via a 512-unit linear layer. All other network components (e.g., actor and critic) remain unchanged from the original implementation.
Training Details and Hyperparameters: We evaluate TD-MPC2 with the wall-time-efficient mode (using 32 parallel GPU environments). Due to the large number of hyperparameters and environments, we do not list them all here. Instead, we refer readers to our public training runs on the Wandb page for TD-MPC2 (state and RGB) results, which include full metrics and configurations.
We only tune two parameters, the replay buffer size and the environment control mode. For all vision-based tasks, we use a smaller replay buffer size of 200K, while for state-based tasks we use 1M. Generally, we employ the delta joint position control mode for wall-time-efficient runs as the environment runs faster relative to IK-based end-effector control; However, for tasks like PickCube-v1, in which the robotic arm cannot easily remain static at the goal, we use the delta end-effector position control mode to achieve better performance. TD-MPC2 baseline also has both a sample-efficient set of environment configurations. The sample-efficient setting uses 1 CPU environment whereas the wall-time efficient setting uses 32 GPU parallelized environments. Less parallel environments relative to the number of updates per frame of data sampled generally lead to better sample efficiency.
X Imitation Learning Baselines
X-A Behavior Cloning Baselines
We evaluate three offline imitation learning baselines on PickCube, PushCube, StackCube, and PegInsertionSide using state and RGB observations. The baselines are Behavior Cloning (BC), Diffusion Policy (DP), and Action Chunking Transformer (ACT). Motion planning is used for the data sources. Table II and III show the best success rates obtained during training on different tasks for state and RGB observations, respectively. Overall we find that Diffusion Policy performs the best, especially when there are few demonstrations.
PickCube-v1 PushCube-v1 StackCube-v1 PegInsertionSide-v1
Demos Success Demos Success Demos Success Demos Success
BC 1000 0.03 1000 0.81 1000 0.00 1000 0.00
ACT 1000 1.00 1000 0.97 1000 0.97 1000 0.43
DP 1000 1.00 1000 0.96 1000 0.99 1000 0.66
BC 100 0.00 100 0.69 100 0.00 100 0.00
ACT 100 0.70 100 0.99 100 0.50 100 0.14
DP 100 1.00 100 0.95 100 0.90 100 0.38
TABLE II: Comparison of imitation learning baselines on different tasks using state observations with varying number of demonstrations.
PickCube-v1 PushCube-v1 StackCube-v1 PegInsertionSide-v1
Demos Success Demos Success Demos Success Demos Success
BC 1000 0.03 1000 0.81 1000 0.00 1000 0.00
ACT 1000 0.98 1000 0.89 1000 0.80 1000 0.00
DP 1000 1.00 1000 0.86 1000 0.81 1000 0.00
BC 100 0.00 100 0.00 100 0.00 100 0.00
ACT 100 0.28 100 0.30 100 0.33 100 0.00
DP 100 0.76 100 0.41 100 0.61 100 0.00
TABLE III: Comparison of imitation learning baselines on different tasks using RGB observations with varying number of demonstrations.
X-B Vision-Language Action (VLA) Model Baselines
We support training and/or evaluation of various vision-language action (VLA) model baselines. For evaluation, we support Octo and RT-x via the GPU parallelized version of the SIMPLER [31] project. For RDT-1B, we support fine-tuning RDT-1B on a few ManiSkill3 demonstrations to then be evaluated on various tasks.
For both training and evaluation, a more classical sense-plan-act style method PerAct[45] is provided as a baseline in ManiSkill3. Different from RGB-based baselines like Octo and RT-x, PerAct operates on voxelized inputs. Our implementation follows the original PerAct design with modifications to demonstration generation and action execution: We obtain demonstrations by replaying trajectories with point cloud observations, which are voxelized directly during inference. This approach eliminates the need to explicitly convert RGBD observations to point clouds, as required in the original implementation. We hard-coded text descriptions for each task based on ManiSkill3 task cards. During inference, the end-effector pose-based actions are converted into joint position sequences using MPlib, which are then executed under ManiSkill3’s joint-position (pd_joint_pos) control mode.
We trained PerAct on PushCube-v1 and StackCube-v1 using 50 demonstration trajectories and evaluated over 100 episodes. The results are presented in Table IV and Figure 27.
Figure 27: Success rate curves of PerAct over 80k training steps on PushCube-v1 and StackCube-v1
X-B1 Multi-View and SE(3) Augmentation
We also reported ablation results under multi-view and SE(3) augmentation settings: The multi-view settings include 4 cameras: a base camera, a left shoulder camera on the left of the arm, a right shoulder camera on the right, and a wrist camera mounted on the arm. Figures 28 and 29 illustrate that multi-view cameras provide finer voxelization details. Yet the difference for StackCube-v1 is marginal since the cubes and their surroundings are already visible via the hand and base cameras. Table V details the cameras used for each experiment.
(a) Default-view voxelized scene.
(b) Multi-view voxelized scene.
Figure 28: Comparison between default-view and multi-view voxelized images in StackCube-v1. The red dot indicates the predicted action.
(a) Default-view RGBD observations. Cameras from left to right: hand_camera, base_camera
(b) Multi-view RGBD observations. From left to right: hand_camera, base_camera, left_shoulder_camera, right_shoulder_camera
Figure 29: Comparison between default-view and multi-view RGBD observations in StackCube-v1. Each column shows one camera’s observations.
Setting PushCube-v1 StackCube-v1
No aug., no m.v. 0.96 0.38
aug., no m.v. 0.98 0.47
aug., m.v. 1.00 0.55
TABLE IV: Best success rates of PerAct over 80k training steps for different tasks and settings. aug.: demo augmentation. m.v.: multi-view
Task Default Cameras Multi-View Cameras
PushCube-v1 base base, l_shldr, r_shldr, wrist
StackCube-v1 base, wrist base, l_shldr, r_shldr, wrist
TABLE V: Camera configurations for each task and setting. l_shldr = left_shoulder, r_shldr = right_shoulder.
The SE(3) augmentation follows the PerAct approach, perturbing point clouds and actions with random translations ([$\pm$0.025m $\pm$0.025m $\pm$0.025m]) and rotations ([$\pm$0 $\pm$0 $\pm$5]).
To summarize the observations and results: (1) SE(3) augmentation helps improve the success rate, though making convergence slower. (2) Multi-view observations can speed up convergence and enhance success rates by providing a more comprehensive 3D scene understanding. (3) As the number of cameras increases, the point cloud size grows. However, voxelized inputs help maintain consistent GPU memory consumption despite the increased data size.
XI Simulation and Rendering Benchmarking